18 Pocuer series Soution (x2+1) y+ xy +zx=0, gCO) = 2, y CO)=3
Answer #1

Power series 16?+o) y + xy+ 2xy =0 , 410)=2 = 2 . y10)=-3 Solution. Given that 122+197 + xy + 20 Dividing by (x! +1), canCnrix (be?+1) I min-1)( xn-2 + x nenxenal + 2x enah n22 nol nio m-2 nen-1) CnX cna + + W nenan a nen-1) cnx n=2 n=2 n=1 + 2ylo)Co = 2 therefor! Co-2 intial condiction given y (0)=2 and y= C, +264 x + 363x2+ y 10) = 4, 5.3 C =3 663 +9, +2 Co= .663C4 ין 2 1 Putting n=3 cs =-(9Cg+26) 20 (9x-1 + 2x0) - B2 20 = 3 2. 20 Cg 17 3 3 40 P of Co, C, ,C2, C3, C4, Cg etc in putting